OpenStudy (anonymous):

Hooke's Law. The distance d when a spring is stretched by a hanging object varies directly as the weight w of the object. If the distance is 17 cm when the weight is 3 kg, what is the distance when the weight is 9 kg?

OpenStudy (anonymous):

varies directly -> y/x = constant hence, distance/weight = constant 17/3 = d/9 d = 17 x 3 = 51